A coffee cup calorimeter is prepared, containing 100.000 g of water (specific heat capacity = 4.184 J/g K) at initial temperature 80.000 C. A salt weighing 5.451 g is quickly added. The salt has a molar mass of 124.742 g/mol. The final temperature of the solution is 73.937 C.

Assume no heat loss to the surroundings. Assume the specific heat capacity of the solution is equal to that of pure water, and that the mass of the solution is equal to the mass of the solid plus the mass of water in the calorimeter.

What is the molar heat of solution for the salt, in kJ/mol?

Report your answer to three digits after the decimal.
Answer #1

mass of the solution = 100+5.451   = 105.451g

no of moles of salt = W/G.M.Wt

                              = 5.451/124.742  

                              = 0.0437 moles

q   = mc\DeltaT

      = 105.451*4.184*(73.937-80)

      = -2675J

         = -2675/0.0437

          = -61213J/mole

         = -61.213KJ/mole >>>>answer